OpenStudy (rockycatsrule):

m to the -5 power = 1/32 HELP WILL FAN AND METAL :)

OpenStudy (rockycatsrule):

Attachments: 24880b0981be7f72c151720ddb9f8ae0a7c3c186.ht

OpenStudy (michele_laino):

hint: we can rewrite suchequation, as below: \[\huge \frac{1}{{{m^5}}} = \frac{1}{{32}}\]

OpenStudy (michele_laino):

oops.. I meant ...such equation...

OpenStudy (michele_laino):

now, please keep in mind that: \[\huge \frac{1}{{32}} = \frac{1}{{{2^5}}}\] since \(32=2^5\), so we can write: \[\huge \frac{1}{{{m^5}}} = \frac{1}{{{2^5}}}\]

OpenStudy (rockycatsrule):

confused

OpenStudy (michele_laino):

we have this property: \[\huge {m^{ - 5}} = \frac{1}{{{m^5}}}\] am I right?

OpenStudy (rockycatsrule):

no

OpenStudy (michele_laino):

why?

OpenStudy (rockycatsrule):